OpenStudy (anonymous):

log[16] 32 = x + 3. explain the procedure used to solve this equation.

OpenStudy (anonymous):

\[\log_{16} 32 = x + 3 \implies \frac 5 4= x+3\implies x=12 \]

OpenStudy (anonymous):

\[\log_{16}32=x+3 \] \[32=16^{x+3}\] \[32=16^{x}*16^3\] \[32/(16^3)=16^{x}\] \[1/128=16^x\] \[\log_{16}(1/128) = x \] \[x = (\log_{10} (1/128))/(\log_{10} (16))\] \[x=-1.75\]

OpenStudy (radar):

This is more realistic (I think), at least the procedure is outlined step by step.

OpenStudy (anonymous):

It checks on substitution as well...12 is not correct.

OpenStudy (radar):

I didn't think so, and the 5/4=x+3 with x=12 was confusing.

OpenStudy (radar):

5/4=x+3 comes out same as your answer -1.75

OpenStudy (anonymous):

Too many steps though, quicker still: \[\log_{16} 32=x+3\] \[\ln {32}/\ln{16}=x+3\] \[x = (\ln32/\ln16) - 3\] \[x = -1.75\]

OpenStudy (anonymous):

@FoolForMath please review your answer as it is marked best and @mape305 may believe it to be correct

OpenStudy (radar):

Just an algebraic "mis-step" he had the correct exponent for 16 to equal 32 (5/4).

OpenStudy (anonymous):

Indeed, just a typo in the end,\[\frac 5 4= x+3\implies 4x=-7 \implies x= -\frac 74\]